OpenStudy (anonymous):

x(sqrt 8/32x) steps please =]

OpenStudy (anonymous):

is it :\[x(\sqrt{8}/32x) \] ? ^_^

OpenStudy (anonymous):

sqrt of the whole inside

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

so we have :\[x(\sqrt{8/32x)}\] let's simplify what's in the sqrt and we'll get : \[x(\sqrt{1/4x})\] you can take 1/4 outside since the sqrt of 4 is 2 and sqrt of 1 is 1 :) : \[=( x/2)(\sqrt{1/x})\] correct me if I'm wrong ^_^

OpenStudy (anonymous):

sorry its the sqrt of 8 divided by sqrt of 32x is that the same thing as your first step?

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

you'll still get the same answer I guess, but all you have to do is find the sqrt of the top and bottom , that's it :)

OpenStudy (anonymous):

ok thanks =]

OpenStudy (anonymous):

np :)